For a linear function f, f(-1)=3 and f(2) = 4.
Find an equation for f.
Find f(3)

Answer by richwmiller(17219)

You have two points
(-1,3) and (2, 4)
the general linear equation in slope intercept form is
y=mx+b
we have
3=-1m+b,
4=2m+b
solve this system of equations for m and b
subtract
-1=-3m
1/3=m
3=-(1/3)+b
3 1/3=b
10/3=b
now we have
y=x/3+10/3
f(3)=3/3+10/3
f(3)=4 1/3 or 13/3